Back to main Archive News page.
Mon 30 Apr 2007 - All systems ran without problems
Fri 27 Apr 2007 - All systems ran without problems
Thu 26 Apr 2007 - All systems ran without problems
Wed 25 Apr 2007 - All systems ran without problems
LUVLE & Related Services Were Unavailable Between 1pm & 5pm
LUVLE, Online Assessments and Module Feedback (Questionmark Perception) were unavailable between 1pm and 5pm to allow for planned essential server maintenance work to be carried out.
Tue 24 Apr 2007 - Network at Risk - Biological Sciences, Physics, County South
During the network at risk period (7am - 9am) upgrades took place in Physics, Biological Sciences, Environmental Science and County South.
Mon 23 Apr 2007 - All systems ran without problems
Fri 20 Apr 2007 - All systems ran without problems
Delayed Lab Opening - Graduate College ISS PC Lab
The Graduate College ISS PC lab was due to re-open today after a major refurbishment but will now re-open next month (date to be announced).
Thu 19 Apr 2007 - All systems ran without problems
News Group Closure
ISS has provided a newsgroup service for a number of years. This service allows Lancaster users to view newsgroup articles, using a local server, and in some cases post to the newsgroup.
Over recent years the internet has evolved and now provides a plethora of ways to view and post to such newsgroups. The server hardware required to support this service is nearing the end of its life and consequently, it is no longer viable to maintain the service. ISS will therefore be closing the news.lancs.ac.uk news service on Thursday 19th April.
This means all users of this service will need to get their news from other outlets, for example http://groups.google.com
Wed 18 Apr 2007 - WebMail service improvements
During the systems at risk period on the 18th of April, the webmail email
service was unavailable whilst work was carried out to move the system database
to a new machine.
Users of POP/IMAP/Wing were unaffected by this work. Work began at 1pm and was completed by 5pm.
Intermittent Service Breaks on Exchange System
Between 1300 and 1500 work was carried out to install essential security patches to all the Exchange servers. There was intermittent short breaks affecting all users through this period.
LUVLE Service Fully Restored
Some LUVLE sites were unavailable earlier in the week due to a hardware error.
The service has been fully restored and all LUVLE sites are now available
Tue 17 Apr 2007 - All systems ran without problems
ResNet unavailable on Southwest Campus
Due to unplanned maintenance on the core switch ResNet was unavailable in the Southwest of campus in student residence between 11am and 12 pm
LUVLE sites for the Management School and Physics were unavailable due to a hardware problem
Mon 16 Apr 2007 - All systems ran without problems
LUVLE Temporarily Unavailable For Management School & Physics
LUVLE sites for the Management School and Physics are currently unavailable due to a hardware problem.
Fri 13 Apr 2007 - All systems ran without problems
Thu 12 Apr 2007 - Central filestore problem
The earlier problem accessing central filestore through the name \\CENTRAL-FILES has now been resolved. Access to central filestore via the name \\LANCS was unaffected by this issue.
Wed 04 Apr 2007 - All systems ran without problems
Central Filestore Maintenance
Personal folders (H drive) and Departmental storage areas were briefly unavailable between 13:00 and 15:00 on Wednesday 4th April to allow for the application of security updates.
New Virus Circulating - User Action Required
A new and dangerous virus is circulating. The new virus infects your PC 'on
opening the message' or when the message is 'reviewed'.
- a Microsoft patch is being pushed by ISS to all campus PCs.
- everyone should run 'Windows Update' on their home PCs.
Tue 03 Apr 2007 - Network maintenance - Replacement of core switches
During the 'Network at Risk' period on the 3rd April the core switch in Music was replaced. This affected Music, Great Hall, CETAD, Language Centre, Bowland North Annexe and County Main. Work begin at 7am and was completed by 9am.
Mon 02 Apr 2007 - All systems ran without problems
Temporary Closure of Furness B48 PC Lab
The Furness B48 ISS PC lab closed at 09:00 am on 2nd April and re-opened at 5pm on Tuesday 3rd April to allow for the fitting of new PCs. All other ISS PC labs remained open.
Sun 01 Apr 2007 - Software upgrade on Central Web Server
On 1st April 2007, ISS upgraded a part of the software that runs on the central web server www.lancs.ac.uk. If you are a web author and you use PHP scripting you may need to make some minor changes, as detailed on the link below. Other users will be unaffected by this change. More Information